The standard deviation of the running time of a sample of 200 fuses was found to be 110 hours. Find a 99% confidence interval for the standard deviation of all such fuses. 


1
Expert's answer
2021-11-03T09:56:25-0400

Degrees of freedom df=n-1=200-1=199.

"99\%CI=(s\sqrt{\frac{n-1}{\chi^2_{0.005}}},s\sqrt{\frac{n-1}{\chi^2_{0.995}}})="

"=(110\sqrt{\frac{200-1}{254.14}},110\sqrt{\frac{200-1}{151.37}})=(97.34,126.12)."